1/12
Looks like no tags are added yet.
Trial
one run of the event
Outcome
any one result from a scenario
sample space
all of the possible outcomes
ex: flipping a coin to get Heads or Tails
Event
verb + outcome flipping a coin to get a H
Probability formula
P(A) = n(a)/n(U)
n(a): how many we want
n(U): all possible options
x’ means not x
complements formula
P(A’) = 1 - P(A)
P(A) + P(A’) = 1
experimental probability
prob based on running an experiment
theoretical probability
prob. based on theory/on paper
relative frequency
number of times the event occurs/number of trials
larger comparable samples generally give reliable or unreliable estimates?
they give more stable estimates
an estimate need not improve after every new trial
how can I improve the estimate
combine the numerators and the denominators of each event/batch
expected number
np (number of trials x probability of the event)
expectation predicts an average count across repetitions
doesn’t need to be a whole number
not guaranteed based on one set of trials